OpenStudy (anonymous):

Brett has been studying a type of bacteria that doubles every month. Originally, there were 5 bacterial cells. He wants to know how many there will be after 42 months. Which equation should he use?

OpenStudy (camerondoherty):

5^42

OpenStudy (anonymous):

A a42 = 5(2)42 - 1 B a42 = 5(2)(42) C a42 = 5(2)42 D a42 = 5(2)(42 - 1)

OpenStudy (anonymous):

@camerondoherty

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

whats the nth term formula for geometric sequence ?

OpenStudy (anonymous):

is that the An=a1+(n-1)d?

ganeshie8 (ganeshie8):

thats for arithmetic sequence, not geometric

OpenStudy (anonymous):

thats the only formula my teacher gave me

OpenStudy (camerondoherty):

\[A _{n}=A _{1} \times r ^{n-1}\] I think...

ganeshie8 (ganeshie8):

yes thats the formula we need to use here :)

OpenStudy (camerondoherty):

Just a hunch i think a = 5

OpenStudy (camerondoherty):

n=42

ganeshie8 (ganeshie8):

yep!

OpenStudy (anonymous):

then whats the answer..

OpenStudy (camerondoherty):

Well look at the equation and look at your options. If a=5 and n=42 which one is it?

OpenStudy (anonymous):

B?

OpenStudy (camerondoherty):

Not quite

OpenStudy (anonymous):

D

OpenStudy (camerondoherty):

look at the equation i showed you again

OpenStudy (anonymous):

its A

OpenStudy (camerondoherty):

Yep!